### v1

@@ 4 Application and numerical experiment
In this section, we apply the bound obtained in Theorem 3.12 to geometric graphs. Geometric graphs model the position in a $d$-dimensional space, so the result can be useful for controlling objects in the real world. This is a generalization of the result shown in [19].
@@
A geometric graph is a connected, undirected, and unweighted graph $\mathcal{G}(V,\mathcal{E})$ such that $V\subseteq\mathcal{Q}$, where $\mathcal{Q}:=[0,l]^{d}\subseteq\mathbb{R}^{d}$ and $|V|=n$. Notice that there are no constraint on $\mathcal{E}$ at first, but the distances between connected pairs of two nodes are used to define parameters.
@@
    the maximum uncovered diameter
@@
There are some relationships among these parameters as follows:
(19) $s\leq r,\ s\leq 2\gamma,\ \rho\leq r,\ \delta\leq\left(\frac{3r}{s}\right)^{d}.$
Inequalities $s\leq r$ and $s\leq 2\gamma$ can be easily checked by definition. The third inequality comes from
$\rho\leq\frac{d_{\mathrm{E}}(u,v)}{d_{\mathcal{G}}(u,v)}\leq\frac{d_{\mathrm{E}}(u,v)}{1}\leq r.$
The last inequality is obtained by comparing volumes of spheres. For a node $u$, let $B\left(u,r+\frac{s}{2}\right)$ be a $d$-dimensional ball centered in $u$ and with radius $r+\frac{s}{2}$. This ball includes a ball $B\left(v,\frac{s}{2}\right)$ for all $v$ which is a neighbor node of $u$. By definition of $s$, balls $B\left(v,\frac{s}{2}\right)$ are not crossing with each other. Therefore, the volume of $B\left(u,r+\frac{s}{2}\right)$ is larger than the sum of the volumes of $B\left(v,\frac{s}{2}\right)$ for all $v$. Then we obtain
$\left(r+\frac{s}{2}\right)^{d}\geq\delta\left(\frac{s}{2}\right)^{d},$
and, by $s\leq r$,
$(3r)^{d}\geq\delta s^{d},$
which leads the last inequality of (19).
@@
Let $P\in\mathbb{R}^{n\times n}$ be a reversible consensus matrix with invariant measure $\bm{\pi}$, associated with a graph $\mathcal{G}=(V,\mathcal{E})$. Assume that all nonzero entries of $P$ belong to the interval $[p_{\min},p_{\max}]$and that $\mathcal{G}$ is a geometric graph with parameters $(s,r,\gamma,\rho)$ and nodes lying in $\mathcal{Q}=[0,l]^{d}$ in which $\gamma<l/4$. Then
@@
$f_{\mathrm{d}}(n)=\begin{cases}n&\textrm{if }d=1,\\ \log n&\textrm{if }d=2,\\ 1&\textrm{if }d\geq 3,\\ \end{cases}$
@@
Let $P\in\mathbb{R}^{n\times n}$ be a consensus matrix with invariant measure $\bm{\pi}$, and let $\mathcal{G}(P)$ be the undirected associated graph of $P$. Assume that all the nonzero entries of $P$ belong to the interval $[p_{\min},p_{\max}]$and that $\mathcal{G}(P)$ is a geometric graph with parameters $(s,r,\gamma,\rho)$ and nodes lying in $\mathcal{Q}=[0,l]^{d}$ in which $\gamma<l/4$. Then
@@
$f_{\mathrm{d}}(n)=\begin{cases}n&\textrm{if }d=1,\\ \log n&\textrm{if }d=2,\\ 1&\textrm{if }d\geq 3,\\ \end{cases}$
@@
Proof 4.3.
Theorem 4.1 in [19] has shown that for reversible $P$,
(20) $c_{l}\bar{\mathcal{R}}(\mathcal{G})\leq J(P)\leq c_{u}\bar{\mathcal{R}}(\mathcal{G})$
with $c_{l}$ and $c_{u}$ dependent on $p_{\min},p_{\max},\delta,\pi_{\max}n,\pi_{\min}n$. Using Theorem 3.12 in this paper, (20) can be generalized to
$J(P)\leq c^{\prime}_{u}\bar{\mathcal{R}}(\mathcal{G}(P)),$
without the assumption of reversibility of $P$, where $c^{\prime}_{u}$ depends on $p_{\min}$,$p_{\max}$,$\delta$,$\pi_{\max}n$, and $\pi_{\min}n$. Moreover, if $P^{*}P=PP^{*}$, then
$J(P)\geq c^{\prime}_{l}\bar{\mathcal{R}}(\mathcal{G}(P)),$
with $c^{\prime}_{l}$ dependent on $p_{\min},p_{\max},\delta,\pi_{\max}n,\pi_{\min}n$. Notice that the lower bound of Theorem 3.12 uses $\delta_{\mathrm{in}}$, but because $\delta_{\mathrm{in}}\leq\delta$, we can define $c^{\prime}_{l}$ as a variable dependent on $\delta$ instead of $\delta_{\mathrm{in}}$.
The rest of the proof is totally analogous to Section 6 in [19].
@@
$s_{n}\geq s,\ r_{n}\leq r,\ \gamma_{n}\leq\gamma,\ \rho_{n}\geq\rho.$
This family of graphs is called a family of geometric graphs with bounded parameters, including regular grids, $R$-disks and Delaunay graphs with appropriate parameter bounds [19]. For a family of geometric graphs which satisfies (4), $s_{n},r_{n},\gamma_{n},\rho_{n}$ and the maximum degree $\delta$ are lower and upper bounded as
[…]
which follows from (19). In addition to the assumption of (4), if the ranges of nonzero entries of $P$ and $n\bm{\pi}$ are bounded, then all parameters determining $k^{\prime}_{1},k^{\prime}_{2},q^{\prime}_{1}$ and $q^{\prime}_{2}$ in Theorem 4.2 are lower and upper bounded. Therefore, we obtain the following corollary.
@@
Let $P_{n}\in\mathbb{R}^{n\times n}$ be a family of consensus matrices with invariant measure $\bm{\pi}_{n}$, and let $\mathcal{G}(P_{n})$ be the undirected associated graph of $P_{n}$. Assume that all nonzero entries of $P_{n}$ belong to the interval $[p_{\min},p_{\max}]$ and all entries of $n\bm{\pi}$ belong to the interval $[\bar{\pi}_{\min},\bar{\pi}_{\max}]$ and that the family of $\mathcal{G}(P_{n})$ is a family of geometric graphs with bounded parameters $(s_{n},r_{n},\gamma_{n},\rho_{n})$ and nodes lying in $\mathcal{Q}=[0,l_{n}]^{d}$ in which $\gamma_{n}<l_{n}/4$. Then
@@
    • We choose $n$ nodes in $[0,l_{n}]^{d}$ in the following way:
@@
        We choose the first node in $[0,l_{n}]^{d}$ according to a uniform distribution.
@@
        We repeat selecting the next node in $[0,l_{n}]^{d}$ according to a uniform distribution independently to the previous nodes. If the Euclidean distance between the picked node and any previous node is less than $s$, then we discard this node. We stop repeating when the number of accepted nodes becomes $n$.
@@
    • For all undirected edges $\{u,v\}$, we change $P_{uv}$ to 00 with probability $p_{\mathrm{d}}$, or $P_{vu}$ to 00 with probability $p_{\mathrm{d}}$, or do nothing to $P_{uv}$ and $P_{vu}$ with the probability $1-2p_{\mathrm{d}}$.
@@
Figure 4: The values of $J(P)/\log n$ in $d=2$ (left panel), and $J(P)/\log n$ in $d=3$ (right panel). The solid lines in both panels are the average of $J(P)$ (divided by $\log n$ in the left panel).
@@
The result is shown in Fig. 4. In the figure of $d=2$, the vertical axis shows $J(P)$ divided by $\log n$. The results suggest that the asymptotic growth of $J(P)$ is bounded as Corollary 4.4. In fact, in small $n$, it seems that there is a decreasing feature in $d=2$ and an increasing feature in $d=3$, but in larger $n$ which satisfies $\gamma<l_{n}/4$, the asymptotic behavior predicted in Corollary 4.4 can be read. In addition, the lower bound for $J(P)$ is not guaranteed since our experiment does not assume $P^{*}P=PP^{*}$, but in the results, the value of $J(P)$ appears not to become much smaller. There may be some lower bound that holds for a random geometric graph with high probability.
@@ Appendix A Calculation of $\gamma_{n}$ and $\rho_{n}$ in the numerical experiment
In the numerical experiment in Section 4, we have to judge whether $\gamma_{n}>\gamma$, and whether $\rho_{n}<\rho$. This condition is checked in the following way.

### v2

@@ 4.3 Application to geometric graphs
In this section, we apply the bounds obtained in Theorem 3.11 to geometric graphs. Geometric graphs model the position in a $d$-dimensional space, so the result can be useful for controlling objects in the real world. This is a generalization of the result shown in [21].
@@
A geometric graph is a connected, undirected and unweighted graph $\mathcal{G}(V,\mathcal{E})$ such that $V\subseteq\mathcal{Q}$, where $\mathcal{Q}:=[0,l]^{d}\subseteq\mathbb{R}^{d}$ and $|V|=n$. Notice that there are no constraint on $\mathcal{E}$ at first, but the distances between connected pairs of two nodes are used to define parameters.
@@
    the maximum uncovered radius
@@
Let $P\in\mathbb{R}^{n\times n}$ be a reversible consensus matrix with invariant measure $\bm{\pi}$, associated with a graph $\mathcal{G}=(V,\mathcal{E})$. Assume that all nonzero entries of $P$ belong to the interval $p_{\min},p_{\max}]$and that $\mathcal{G}$ is a geometric graph with parameters $(s,r,\gamma,\rho)$ and nodes lying in $\mathcal{Q}=[0,l]^{d}$ in which $\gamma<l/4$. Then
@@
[…]
@@
Let $P\in\mathbb{R}^{n\times n}$ be a consensus matrix with invariant measure $\bm{\pi}$, and let $\mathcal{G}(P)$ be the associated undirected graph of $P$. Assume that all the nonzero entries of $P$ belong to the interval $p_{\min},p_{\max}]$and that $\mathcal{G}(P)$ is a geometric graph with parameters $(s,r,\gamma,\rho)$ and nodes lying in $\mathcal{Q}=[0,l]^{d}$ in which $\gamma<l/4$. Then
@@
[…]
@@
We omit the proof because it is analogous to Section 6 in [21] except the bounds of $J(P)$ should be evaluated by Theorem 3.11.
@@
(19) $s_{n}\geq s,\ r_{n}\leq r,\ \gamma_{n}\leq\gamma,\ \rho_{n}\geq\rho.$
This family of graphs is called a family of geometric graphs with bounded parameters, including grid graphs with boundaries where the edges are connected only to the nearest neighbors [21]. In this sense, this family can be understood as a generalization of the grid graph case.
Although Proposition 4.6 discuss the case of fixed geometric parameters, the similar result to Proposition 4.6 is also known for this family of geometric graphs with bounded parameters in [21]. Here, we state the generalized version of it as a corollary.
@@
Let $P_{n}\in\mathbb{R}^{n\times n}$ be a family of consensus matrices with invariant measure $\bm{\pi}_{n}$, and let $\mathcal{G}(P_{n})$ be the associated undirected graph of $P_{n}$. Assume that all nonzero entries of $P_{n}$ belong to the interval $p_{\min},p_{\max}]$ and all entries of $n\bm{\pi}$ belong to the interval $\bar{\pi}_{\min},\bar{\pi}_{\max}]$ and that the family of $\mathcal{G}(P_{n})$ is a family of geometric graphs with bounded parameters $(s_{n},r_{n},\gamma_{n},\rho_{n})$ and nodes lying in $\mathcal{Q}=[0,l_{n}]^{d}$ in which $\gamma_{n}<l_{n}/4$. Then
@@
The proof can be done by combining the discussion of Theorem 4.7 and Section 4.3 in [21].
@@
    • We choose $n$ nodes in $0,l_{n}]^{d}$ in the following way:
@@
        We choose the first node in $0,l_{n}]^{d}$ according to a uniform distribution.
@@
        We repeat selecting the next node in $0,l_{n}]^{d}$ according to a uniform distribution independently to the previous nodes. If the Euclidean distance between the picked node and any previous node is less than $s$, then we discard this node. We stop repeating when the number of accepted nodes becomes $n$.
@@
    • For all undirected edges $\{u,v\}$, we change $P_{uv}$ to 0 with probability $p_{\mathrm{d}}$, or $P_{vu}$ to 0 with probability $p_{\mathrm{d}}$, or do nothing to $P_{uv}$ and $P_{vu}$ with the probability $1-2p_{\mathrm{d}}$.
@@
Figure 10: The values of $J(P)/\log n$ in $d=2$ (left panel), and $J(P)$ in $d=3$ (right panel). The solid lines in both panels are the average of $J(P)$ (divided by $\log n$ in the left panel).
@@
The result is shown in Figure 10. In the figure of $d=2$, the vertical axis shows $J(P)$ divided by $\log n$. The results suggest that the asymptotic growth of $J(P)$ is bounded as Corollary 4.8. In fact, in small $n$, it seems that there is a decreasing feature in $d=2$ and an increasing feature in $d=3$, but in larger $n$ which satisfies $\gamma<l_{n}/4$, the asymptotic behavior predicted in Corollary 4.8 can be read. In addition, the lower bound for $J(P)$ is not guaranteed since our experiment does not assume $P^{*}P=PP^{*}$, but in the results, the value of $J(P)$ appears not to become much smaller. There may be some lower bound that holds for a random geometric graph with high probability.
@@ Appendix A Calculation of $\gamma_{n}$ and $\rho_{n}$ in the numerical experiment
In the numerical experiment in Section 4.3, we have to judge whether $\gamma_{n}>\gamma$, and whether $\rho_{n}<\rho$. This condition is checked in the following way.